an angle is 14 degrees more than the measure of its complement. find the number of degrees in each angle
Answer by rapaljer(4551)

Let x = first angle
x+14= second angle

x+x+14 = 90
2x +14= 90
2x=76
x= 38 degrees
x+14= 52 degrees

Check: 38+52 = 90 degrees

R^2